# Pairing by disorder of dopants in a magnetic spin ladder

### v1

@@ Abstract
Understanding pairing of dopants in magnetic lattices is an immense challenge in modern quantum many-body theory. The main goal is to gain a deeper understanding of strongly correlated systems, arising e.g. in the notoriously difficult high-temperature superconductors Schrieffer and Brooks (2007); Trugman (1988); Dagotto (1994). While solid state experiments continue to pursue such inquiries after decades of research O’Mahony et al. (2022), a highly interesting and promising pathway has appeared in quantum simulation experiments with ultracold atoms in optical lattices Gross and Bloch (2017); Ji et al. (2021); Wang et al. (2021); Prichard et al. (2023); Koepsell et al. (2019, 2021); Lebrat et al. (2024); Hirthe et al. (2023). Single-site resolution capabilities Bakr et al. (2009); Sherson et al. (2010) makes it possible to measure low-temperature spatial correlations Koepsell et al. (2019, 2021); Prichard et al. (2023); Lebrat et al. (2024), as well as track them over time by repeated realizations of the same initial state. In particular, single hole dynamics consistent with magnetic polaron formation Nielsen et al. (2022) in quantum antiferromagnets has been measured Ji et al. (2021). Moreover, in such experiments one can tinker with the simulated models, otherwise fixed in the solid state. This has enabled the observation of dopant pairing in a mixed-dimensional magnetic lattice Hirthe et al. (2023). This has also been seen to be a plausible explanation Schlömer et al. (2023) for high-temperature superconductivity in a specific nickelate compound Sun et al. (2023). More broadly speaking, the appearance of pairing is conventionally tied to the distortion of some underlying order. In the aforementioned example, it is the distortion of the underlying AFM order that gives rise to the induced attraction Bohrdt et al. (2022); Nielsen (2023), which is then further enhanced by restricting motion to one dimension. In conventional superconductors, electrons distort the underlying crystal lattice, inducing an attraction between them Fröhlich (1950); Bardeen et al. (1957); Bruus and Flensberg (2016). In materials supporting high-temperature superconductivity, there may be an important counterexample to this generic scenario, in which preformed Cooper pairs Geshkenbein et al. (1997); Randeria et al. (1992) appearing above the superconducting phase transition might explain Li et al. (2010); Seo et al. (2019) the pseudogap phase Stanescu and Phillips (2003).
@@
Figure 1: Pairing dynamics. (a) Setup: holes move in a magnetic (red: \!\!$\uparrow$, blue: \!\!$\downarrow$) two-leg ladder with Ising interactions $J_{\perp},J_{\parallel}$ via nearest-neighbor hopping $t$ along the ladder. (b) Hole-hole correlator $C_{\rm hh}(x_{1},x_{2};\tau)=\braket{\hat{n}_{h}(x_{1})\hat{n}_{h}(x_{2})}_{\tau}$ for $J_{\parallel}=0$ at indicated times normalized to its maximum after the holes are initialized on the same rung, $(l,x_{1})=(1,0),(l,x_{2})=(2,0)$. (c) Dynamics of the root-mean-square dynamics of the center-of-mass $X=(x_{1}+x_{2})/2$ and relative $\Delta=(x_{1}-x_{2})/2$ coordinates, (d) correlation coefficient ${\rm corr}(x_{1},x_{2})=\braket{x_{1}x_{2}}_{\tau}/[\sigma(x_{1})\sigma(x_{2})]$, and (e) relative correlation function for $\Delta=0$. Steady state values are shown in short dashed lines. In all: $J_{\perp}=5t$.
@@
In this Letter, we demonstrate a pairing mechanism that does not rely on being in nor in the vicinity of a phase transition to an ordered phase. In particular, the discovered mechanism remains at infinite temperatures and is a direct consequence of the thermally disordered spin lattice. I analyze the effect in a simplistic setup of a mixed-dimensional model [Fig. 1(a)], where two holes are placed in two legs of a spin ladder. The spin degree of freedom is assumed to be at infinite temperature and completely disordered. Conventional wisdom, therefore, suggests that the holes should not be able to pair. Nevertheless, when the holes are initialized on the same rung in the infinite temperature spin lattice, the strong thermal disorder of the system frustrates the relative motion of the dopants. Indeed, tracking the hole-hole correlator in time [Fig. 1(b)], shows a significant correlation along the diagonal $x_{1}=x_{2}$, where the holes are on the same rung of the ladder. One may equivalently observe that the inter-hole distance is always smaller than the traversed distance of their center-of-mass [Fig. 1(c)]. This results in a strictly positive correlation between the holes’ position, signaling an effective attraction between them [Fig. 1(d)]. Finally, we observe that the probability of finding the holes on the same rung is easily increased by a factor of $4$ as compared to what one expects from their marginal distributions, described by the relative hole-hole correlator in Fig. 1(e). This phenomenon is closely tied to the recent discovery of thermally induced localization in such systems Nielsen (2024a, b). In particular, the holes experience an effective induced interaction that has a disordered character, encouraging the holes to co-propagate. I emphasize that the discovered effect would appear in a wide range of systems, though the robustness of the pairing in more generic setups is presently unclear.
@@
i.e. a mixed-dimensional $t$-$J_{z}$ model Grusdt and Pollet (2020). I focus on the case of nearest-neighbor couplings along, $J_{\parallel}=J_{\pm{\bf e}_{x}}$, and across, $J_{\perp}=J_{\pm{\bf e}_{y}}$, the ladder, with $\hat{S}^{(z)}_{\bf i}$ the $z$-projection of the spin operator $\hat{\bf S}_{\bf i}$. The hopping, ${\hat{H}}_{t}$, is only allowed along the ladder and only to vacant sites, enforced by the constrained operators $\tilde{c}^{\dagger}_{\sigma{\bf i}}=\hat{c}^{\dagger}_{\sigma{\bf i}}[1-\hat{n}_{\bf i}]$, where $\hat{n}_{\bf i}=\sum_{\sigma}\hat{n}_{\sigma{\bf i}}=\sum_{\sigma}\hat{c}^{\dagger}_{\sigma{\bf i}}\hat{c}_{\sigma{\bf i}}$ is the local density operator. The effective spin degree of freedom generally describes an internal degree of freedom with the Ising-type coupling in ${\hat{H}}_{J}$, the first term in Eq. (1), and since there are no exchange processes of the dopants, both fermions and bosons can realize the model. To have an efficient description of the hole and spin degrees of freedom, we perform a Holstein-Primakoff transformation on top of the ferromagnetic ground state $|\ket{{\rm FM}}=\ket{\cdots\uparrow\uparrow\cdots}$, with all spins in the $\ket{\uparrow}$ state. The spin Hamiltonian
@@
hereby, describes the interaction of bosonic spin flip excitations $\hat{s}^{\dagger}_{\bf i}$, corresponding to the creation of a spin-$\downarrow$ on site ${\bf i}$, and holes created by the operator $\hat{h}^{\dagger}_{\bf i}$, which inherit the statistics of the underlying particles, be it fermionic or bosonic Nielsen (2023). Equivalently, the hopping Hamiltonian
@@
describes two distinct ways in which holes may hop: (1) a hole hopping from site ${\bf i}$ to ${\bf j}$ in the absence of a spin flip on site ${\bf j}$, and (2) hopping in the presence of a spin flip, whereby the hole and spin flip swap places. The constraint function $F(\hat{h},\hat{s})=\sqrt{1-\hat{s}^{\dagger}\hat{s}-\hat{h}^{\dagger}\hat{h}}$ ensures at most one spin per site. The initial state at time $\tau=0$ of two holes on the same rung of the ladder in a spin background ${\boldsymbol{\sigma}}$ is then
|$\ket{\Psi_{\boldsymbol{\sigma}}(\tau=0)}=\prod_{l=1}^{2}\hat{h}^{\dagger}_{(l,0)}\prod_{j\in S_{{\boldsymbol{\sigma}}}^{l}}\hat{s}^{\dagger}_{(l,j)}\ket{{\rm FM}},$ (4)
in which $S_{{\boldsymbol{\sigma}}}^{l}$ is the subset of spins in leg $l$, which are initially in the $\ket{\downarrow}$ state. Here, we write ${\bf i}=(l,j)$, with leg index $l=1,2$ and position $j$ along the leg. Since the spins can only exchange places with the holes along the ladder, the state at any later stage is
@@
Here, ${\rm sgn}(x)$ is the sign function. Despite the rather formal description, the physics is simple and clear: if the hole in leg $l$ moves $|x_{l}|$ times to the right, it surpasses $|x_{l}|$ spins that all move one step to the left. If the hole moves to the left, the surpassed spins move one step to the right. The corresponding probability amplitudes $a_{{\boldsymbol{\sigma}}}(x_{1},x_{2};\tau)$ of finding the holes at positions $x_{1},x_{2}$ at time $\tau$, follow the equations of motion
@@
Figure 2: Pairing mechanism. (a) As the holes move (green shaded region), spin bonds in the initial configuration are broken across (dark blue) and within (light blue) the ladder, and new bonds are established (dark and light red, respectively), resulting in a random induced interaction between the holes. (b) Induced interaction for vanishing center-of-mass, $X=0$, as a function of the relative coordinate $\Delta$ for $50$ realizations of the spin background (grey and colored lines), performing a random walk with standard deviation $\sim|J_{\perp}|\sqrt{|\Delta|}$ (black lines). (c) Induced interaction at $\Delta=0$, as a function of $X$. Here, $J_{\parallel}=J_{\perp}$ is used.
I indeed show SM that if all spin backgrounds are equally likely – as in an infinite temperature spin state – $V_{\boldsymbol{\sigma}}(x_{1},x_{2})$ has zero mean and a variance ($x_{1},x_{2}\neq 0$)
${\rm Var}[V_{{\boldsymbol{\sigma}}}(x_{1},x_{2})]=\frac{J_{\perp}^{2}}{4}\left[|\Delta|-\frac{1}{4}\right]+\frac{J_{\parallel}^{2}}{4}$ (7)
@@
linear in the relative distance $2|\Delta|$ between the holes, describing a classical random walk in the values of the potential [Fig. 2(b)]. I can focus on the evolution in specific spin backgrounds by resolving the dynamics of the density matrix in terms of these states. Starting from the initial density matrix
$\hat{\rho}(\tau=0)=\!\!\!\sum_{l=1,2,\sigma_{(l,0)}}\!\!\!\hat{c}_{(l,0)\sigma{(l,0)}}\hat{\rho}_{J}\hat{c}^{\dagger}_{(l,0)\sigma{(l,0)}},$ (8)
where $\hat{\rho}_{J}=e^{-\beta{\hat{H}}_{J}}/Z_{J}$ is the Gibbs state of the spins at inverse temperature $\beta=1/(k_{B}T)$ and $\sigma_{(l,0)}=\uparrow,\downarrow$ designates the spin configuration at the origin of each leg, $l$, the ensuing dynamics $\hat{\rho}(\tau)=e^{-i\hat{H}\tau}\hat{\rho}(\tau=0)e^{+i\hat{H}\tau}$ – provided that the system is assumed to be closed – is the Boltzmann-weighted sum Nielsen (2024a, b)
$\hat{\rho}(\tau)=\sum_{{\boldsymbol{\sigma}}}\frac{e^{-\beta E_{J}({\boldsymbol{\sigma}})}}{Z_{J}}\ket{\Psi_{{\boldsymbol{\sigma}}}(\tau)}\bra{\Psi_{{\boldsymbol{\sigma}}}(\tau)},$ (9)
@@
where $E_{J}({\boldsymbol{\sigma}})$ is the magnetic energy of the spin realization ${\boldsymbol{\sigma}}$ before the spins $\sigma_{(1,0)},\sigma_{(2,0)}$ are removed. In particular, at infinite temperatures all pure state evolutions |$\ket{\Psi_{{\boldsymbol{\sigma}}}(\tau)}$ are sampled with equal probability, $e^{-\beta E_{J}({\boldsymbol{\sigma}})}/Z_{J}=2^{-N}$, for system size $N$. In this high-temperature limit, the dynamics is exactly the same for antiferromagnetic ($J_{\perp},J_{\parallel}>0$) and ferromagnetic couplings ($J_{\perp},J_{\parallel}<0$). The calculation of the hole dynamics then follows the recipe: (1) generate a sample $S$ of $N_{S}=400$ random spin configurations ${\boldsymbol{\sigma}}$ and calculate $V_{{\boldsymbol{\sigma}}}(x_{1},x_{2})$, (2) solve the equations of motion in Eq. (6) using exact diagonalization for $N=161\times 2$, and (3) compute the hole-hole correlator
$C_{\rm hh}(x_{1},x_{2};\tau)=\braket{\hat{n}_{h}(x_{1})\hat{n}_{h}(x_{2})}_{\tau}=\sum_{{\boldsymbol{\sigma}}\in S}\frac{|a_{{\boldsymbol{\sigma}}}(x_{1},x_{2};\tau)|^{2}}{N_{S}},$ (10)
@@
whereby it may be tracked in time [Fig. 1(b)]. This describes the probability of simultaneously finding holes at positions $(1,x_{1}),(2,x_{2})$. Here, $\hat{n}_{h}(x_{l})=\hat{h}^{\dagger}_{(l,x_{l})}\hat{h}_{(l,x_{l})}$ is the hole density operator. Figure 1(b) manifestly breaks the individual inversion symmetries $x_{1}\to-x_{1}$ or $x_{2}\to-x_{2}$ of uncorrelated holes, and shows a clear tendency for them to be on the same rung ($x_{1}=x_{2}\Leftrightarrow\Delta=0$). Their motion is further analyzed by the root-mean-square of the center-of-mass and relative coordinates [Fig. 1(c)],
$A_{\rm rms}(\tau)$ $=\left[\sum_{X,\Delta}A^{2}C_{\rm hh}(X+\Delta,X-\Delta;\tau)\right]^{1/2},$ (11)
@@
which is always restricted to the interval $[-1,+1]$. Positive values of ${\rm corr}(x_{1},x_{2})$, as seen in Fig. 1(d), correspond to $X_{\rm rms}>\Delta_{\rm rms}$, and signify attraction. Likewise, negative values correspond to an effective repulsion, while ${\rm corr}(x_{1},x_{2})=0$ corresponds to no linear correlations. To further substantiate the findings, I analyze the relative hole-hole correlator
$C_{\rm hh}^{\rm rel}(\Delta,\tau)=\frac{\sum_{X}C_{\rm hh}(X+\Delta,X-\Delta;\tau)}{\sum_{X}C_{\rm hh}^{\rm disc}(X+\Delta,X-\Delta;\tau)},$ (13)
giving the probability of finding the holes at distance $2|\Delta|$ relative to what is expected from their marginals, i.e. the disconnected correlator $C^{\rm disc}_{\rm hh}(\Delta,\tau)=\sum_{X}\braket{\hat{n}_{h}(X+\Delta)}\braket{\hat{n}_{h}(X-\Delta)}$. This is shown for $\Delta=0$ in Fig. 1(e), demonstrating a significant increase in the probability of finding the holes on the same rung. On long timescales, the system settles into a steady state, by which $X_{\rm rms}\to l_{X},\Delta_{\rm rms}\to l_{\Delta}$. For $J_{\perp}\geq 2t$, there are negligible finite-size corrections for the chosen length of $N/2=161$, and I, therefore, confine the analysis to this regime. Figure 3(a), hereby, shows that $l_{X}>l_{\Delta}$ remains true over this entire range of parameters. The associated asymptote of the correlation coefficient [Fig. 3(b)] shows very strong attractive correlations for $J_{\perp}\gg t$ and $J_{\parallel}=0$.

### v3

# Dopant pairing in a disordered magnetic spin ladder

@@ Abstract
Understanding pairing of dopants in magnetic lattices is an immense challenge in modern quantum many-body theory. The main goal is to gain a deeper understanding of strongly correlated systems, arising e.g. in the notoriously difficult high-temperature superconductors [1, 2, 3]. While solid state experiments continue to pursue such inquiries after decades of research [4], a highly interesting and promising pathway has appeared in quantum simulation experiments with ultracold atoms in optical lattices [5, 6, 7, 8, 9, 10, 11, 12]. Single-site resolution capabilities [13, 14] makes it possible to measure low-temperature spatial correlations [9, 10, 8, 11], as well as track them over time by repeated realizations of the same initial state. In particular, single hole dynamics consistent with magnetic polaron formation [15] in quantum antiferromagnets has been measured [6]. Moreover, in such experiments one can tinker with the simulated models, otherwise fixed in the solid state. This has enabled the observation of dopant pairing in a mixed-dimensional magnetic lattice [12], also seen to be a plausible explanation [16] for high-temperature superconductivity in a specific nickelate compound [17]. More broadly speaking, the appearance of pairing is conventionally tied to the distortion of some underlying order, be it the underlying antiferromagnetic order in the above example [18, 19], or how, in conventional superconductors, electrons distort the underlying crystal lattice, inducing an attraction between them [20, 21, 22]. In materials supporting high-temperature superconductivity, there may be an important counterexample to this generic scenario, in which preformed Cooper pairs [23, 24] appearing above the superconducting phase transition might explain [25, 26] the pseudogap phase [27].
@@
Figure 1: Pairing dynamics. (a) Setup: holes move in a magnetic (red: \!\!$\uparrow$, blue: \!\!$\downarrow$) two-leg ladder with Ising interactions $J_{\perp},J_{\parallel}$ via nearest-neighbor hopping $t$ along the ladder. (b) Hole-hole correlator $C_{hh}(x_{1},x_{2};\tau)=\braket{\hat{n}_{h}(x_{1})\hat{n}_{h}(x_{2})}_{\tau}$ for $J_{\parallel}=0$ at indicated times normalized to its maximum after the holes are initialized on the same rung, $(l,x_{1})=(1,0),(l,x_{2})=(2,0)$. (c) Dynamics of the root-mean-square dynamics of the center-of-mass $X=(x_{1}+x_{2})/2$ and relative $\Delta=(x_{1}-x_{2})/2$ coordinates, (d) correlation coefficient ${\rm corr}(x_{1},x_{2})=\braket{x_{1}x_{2}}_{\tau}/[\sigma(x_{1})\sigma(x_{2})]$, and (e) relative correlation function for $\Delta=0$. Steady state values are shown in short dashed lines. In all: $J_{\perp}=5t$.
@@
In this Letter, I demonstrate a pairing mechanism that does not rely on being in nor in the vicinity of a phase transition to an ordered phase. In particular, the discovered mechanism remains at infinite temperatures and is a direct consequence of the thermally disordered spin lattice. I analyze the effect in a simplistic setup of a mixed-dimensional model [Fig. 1(a)], where two holes are placed in two legs of a spin ladder. The spin degree of freedom is assumed to be at infinite temperature and completely disordered. Conventional wisdom, therefore, suggests that the holes should not be able to pair. Nevertheless, when the holes are initialized on the same rung in the infinite temperature spin lattice, the strong thermal disorder of the system frustrates the relative motion of the dopants. Indeed, tracking the hole-hole correlator in time [Fig. 1(b)], shows a significant correlation along the diagonal $x_{1}=x_{2}$, where the holes are on the same rung of the ladder. One may equivalently observe that the inter-hole distance is always smaller than the traversed distance of their center-of-mass [Fig. 1(c)]. This results in a strictly positive correlation between the holes’ positions, signaling an effective attraction between them [Fig. 1(d)]. Finally, we observe that the probability of finding the holes on the same rung is easily increased by a factor of $4$, as compared to what one expects from their marginal distributions. See the relative hole-hole correlator in Fig. 1(e). This phenomenon arises due to an effective induced interaction that has a disordered character, encouraging the holes to co-propagate, and is closely tied to the recent discovery of thermally induced localization in such systems [28, 29]. In the extreme limit of $J_{\parallel}=0$, the holes even dislodge from their original positions and spread subdiffusively as a pair. I emphasize that the discovered effect would appear in a wide range of systems, though the robustness of the pairing in more generic setups is presently unclear.
@@
i.e. a mixed-dimensional $t$-$J_{z}$ model [30]. I focus on the case of nearest-neighbor couplings along, $J_{\parallel}=J_{\pm{\bf e}_{x}}$, and across, $J_{\perp}=J_{\pm{\bf e}_{y}}$, the ladder, with $\hat{S}^{(z)}_{\bf i}$ the $z$-projection of the spin operator $\hat{\bf S}_{\bf i}$. The hopping, ${\hat{H}}_{t}$, is only allowed along the ladder and only to vacant sites, enforced by the constrained operators $\tilde{c}^{\dagger}_{\sigma{\bf i}}=\hat{c}^{\dagger}_{\sigma{\bf i}}[1-\hat{n}_{\bf i}]$, where $\hat{n}_{\bf i}=\sum_{\sigma}\hat{n}_{\sigma{\bf i}}=\sum_{\sigma}\hat{c}^{\dagger}_{\sigma{\bf i}}\hat{c}_{\sigma{\bf i}}$ is the local density operator. The effective spin degree of freedom generally describes an internal degree of freedom with the Ising-type coupling in ${\hat{H}}_{J}$, the first term in Eq. (1), and since there are no exchange processes of the dopants, both fermions and bosons can realize the model. To have an efficient description of the hole and spin degrees of freedom, I perform a Holstein-Primakoff transformation on top of the ferromagnetic ground state $|\ket{{\rm FM}}=\ket{\cdots\uparrow\uparrow\cdots}$, with all spins in the $\ket{\uparrow}$ state. The spin Hamiltonian
@@
hereby, describes the interaction of bosonic spin flip excitations $\hat{s}^{\dagger}_{\bf i}$, corresponding to the creation of a spin-$\downarrow$ on site ${\bf i}$, and holes created by the operator $\hat{h}^{\dagger}_{\bf i}$, which inherit the statistics of the underlying particles, be it fermionic or bosonic [19]. Equivalently, the hopping Hamiltonian
@@
describes two distinct ways in which holes may hop: (1) a hole hopping from site ${\bf i}$ to ${\bf j}$ in the absence of a spin flip on site ${\bf j}$, and (2) hopping in the presence of a spin flip, whereby the hole and spin flip swap places. The constraint $F(\hat{h},\hat{s})=\sqrt{1-\hat{s}^{\dagger}\hat{s}-\hat{h}^{\dagger}\hat{h}}$ ensures at most one spin per site. The initial state at time $\tau=0$ of two holes on the same rung of the ladder in a spin background ${\boldsymbol{\sigma}}$ is then |$\ket{\Psi_{\boldsymbol{\sigma}}(\tau=0)}=\prod_{l=1}^{2}\hat{h}^{\dagger}_{(l,0)}\prod_{j\in S_{{\boldsymbol{\sigma}}}^{l}}\hat{s}^{\dagger}_{(l,j)}\ket{{\rm FM}}$, in which $S_{{\boldsymbol{\sigma}}}^{l}$ is the subset of spins in leg $l$, which are initially in the $\ket{\downarrow}$ state. Here, I write ${\bf i}=(l,j)$, with leg index $l=1,2$ and position $j$ along the leg. Since spins only exchange places with the holes along the ladder, the state at any later stage is
@@
Here, ${\rm sgn}(x)$ is the sign function. Despite the rather formal description, the physics is simple and clear: if the hole in leg $l$ moves $|x_{l}|$ times to the right (left), it surpasses $|x_{l}|$ spins that all move one step to the left (right). The corresponding probability amplitudes $a_{{\boldsymbol{\sigma}}}(x_{1},x_{2};\tau)$ of finding the holes at positions $x_{1},x_{2}$ at time $\tau$ obey
@@
Figure 2: Pairing mechanism. (a) As the holes move (green shaded region), spin bonds in the initial configuration are broken across (dark blue) and within (light blue) the ladder, and new bonds are established (dark and light red, respectively), giving a random induced interaction between the holes. (b) Induced interaction for vanishing center-of-mass, $X=0$, as a function of the relative coordinate $\Delta$ for $50$ realizations of the spin background (grey and colored lines), performing a random walk with standard deviation $\sim|J_{\perp}|\sqrt{|\Delta|}$ (black lines). (c) Induced interaction at $\Delta=0$, as a function of $X$.
@@
I indeed show [31] that if all spin backgrounds are equally likely – as in an infinite temperature spin state – $V_{\boldsymbol{\sigma}}(x_{1},x_{2})$ has zero mean and a variance ${\rm Var}[V_{{\boldsymbol{\sigma}}}(x_{1},x_{2})]\propto J_{\perp}^{2}|\Delta|$, linear in the relative distance $2|\Delta|$ between the holes, describing a classical random walk of $V$ in $|\Delta|$ [Fig. 2(b)]. I then focus on the evolution in specific spin backgrounds by resolving the dynamics of the density matrix in terms of these states. Starting from the initial density matrix
$\hat{\rho}(0)=\!\sum_{\sigma_{1},\sigma_{2}}\hat{c}_{(2,0),\sigma_{2}}\hat{c}_{(1,0),\sigma_{1}}\hat{\rho}_{J}\hat{c}_{(1,0),\sigma_{1}}^{\dagger}\hat{c}_{(2,0),\sigma_{2}}^{\dagger},$ (6)
@@
where $\hat{\rho}_{J}=e^{-\beta{\hat{H}}_{J}}/Z_{J}$ is the Gibbs state of the spins at inverse temperature $\beta=1/(k_{B}T)$ and $\sigma_{l}=\uparrow,\downarrow$ designates the spin configuration at the origin of each leg, $l$, the ensuing dynamics $\hat{\rho}(\tau)=e^{-i\hat{H}\tau}\hat{\rho}(0)e^{+i\hat{H}\tau}$ – provided that the system is assumed to be closed – is the Boltzmann-weighted sum $\hat{\rho}(\tau)=Z_{J}^{-1}\sum_{{\boldsymbol{\sigma}}}e^{-\beta E_{J}({\boldsymbol{\sigma}})}\ket{\Psi_{{\boldsymbol{\sigma}}}(\tau)}\bra{\Psi_{{\boldsymbol{\sigma}}}(\tau)}$ | [28, 29], where $E_{J}({\boldsymbol{\sigma}})$ is the magnetic energy of the spin realization ${\boldsymbol{\sigma}}$ before the spins $\sigma_{1},\sigma_{2}$ are removed. At infinite temperatures all pure state evolutions |$\ket{\Psi_{{\boldsymbol{\sigma}}}(\tau)}$ are sampled equally, $e^{-\beta E_{J}({\boldsymbol{\sigma}})}/Z_{J}=2^{-N}$, for system size $N$, and the ensuing dynamics is exactly equal for antiferromagnetic ($J_{\perp},J_{\parallel}>0$) and ferromagnetic couplings ($J_{\perp},J_{\parallel}<0$). The calculation of the hole dynamics then follows the recipe: (1) generate a sample $S$ of $N_{S}=400$ random spin configurations ${\boldsymbol{\sigma}}$ and calculate $V_{{\boldsymbol{\sigma}}}(x_{1},x_{2})$, (2) solve the equations of motion in Eq. (5) using exact diagonalization for $N=201\times 2$, and (3) compute the hole-hole correlator
$C_{hh}(x_{1},x_{2};\tau)=\braket{\hat{n}_{h}(x_{1})\hat{n}_{h}(x_{2})}_{\tau}=\!\sum_{{\boldsymbol{\sigma}}\in S}\!\frac{|a_{{\boldsymbol{\sigma}}}(x_{1},x_{2};\tau)|^{2}}{N_{S}},$ (7)
@@
as a function of time [Fig. 1(b)]. This describes the probability of simultaneously finding holes at positions $(1,x_{1}),(2,x_{2})$, with $\hat{n}_{h}(x_{l})=\hat{h}^{\dagger}_{(l,x_{l})}\hat{h}_{(l,x_{l})}$ the hole density operator. Figure 1(b) manifestly breaks the individual inversion symmetries $x_{1}\to-x_{1}$ or $x_{2}\to-x_{2}$ of uncorrelated holes, and shows a clear tendency for them to be on the same rung ($x_{1}=x_{2}\Leftrightarrow\Delta=0$). Their motion is further analyzed by the root-mean-square of the center-of-mass and relative coordinates [Fig. 1(c)],
$A_{\rm rms}(\tau)=\left[\sum_{X,\Delta}A^{2}C_{hh}(X+\Delta,X-\Delta;\tau)\right]^{1/2},$ (8)
@@
taking values in $[-1,+1]$. Positive values, as in Fig. 1(d), correspond to $X_{\rm rms}>\Delta_{\rm rms}$, and signify attraction. Likewise, negative values correspond to effective repulsion, while ${\rm corr}(x_{1},x_{2})=0$ corresponds to no linear correlations. To further substantiate the findings, I analyze the relative hole-hole correlator
$C_{hh}^{\rm rel}(\Delta,\tau)=\frac{\sum_{X}C_{hh}(X+\Delta,X-\Delta;\tau)}{C_{hh}^{\rm disc}(\Delta,\tau)},$ (10)
giving the probability of finding the holes at distance $2|\Delta|$ relative to what is expected from their marginals, i.e. the disconnected correlator $C^{\rm disc}_{hh}(\Delta,\tau)=\sum_{X}\braket{\hat{n}_{h}(X+\Delta)}\braket{\hat{n}_{h}(X-\Delta)}$. This is shown for $\Delta=0$ in Fig. 1(e), demonstrating a dynamically increasing probability of finding the holes on the same rung.
For vanishing intra-leg spin coupling, $J_{\parallel}=0$, the holes will continue to spread into the system. Indeed, by analyzing the center-of-mass dynamics [black-filled colored lines in Fig. 3(a)], I find that it transitions from an initial ballistic expansion with velocity $t$ to subdiffusive behavior, $X_{\rm rms}\sim(\tau\cdot t)^{\nu}$. This persistent spreading comes from a vanishing potential for holes on the same rung, $V(x,x)=0$. In fact, in the limit of $J_{\perp}\gg t$, I gain simplified quantitative insights by assuming the holes move at most one site away from each other. The resulting effective model allows me to push the numerics to larger systems and longer times, verifying the subdiffusive behavior, and is seen to perfectly match the full dynamics for large $J_{\perp}/t$ on all timescales [gray-filled colored lines in Fig. 3(a)]. The subdiffusive exponent [bottom of Fig. 3(a)] slowly decreases for increasing $J_{\perp}/t$, reaching a plateau at $\nu\simeq 0.32$ for $J_{\perp}/t\gg 1$.